OpenStudy (anonymous):

Solve using synthetic division: (x^2-3x+2) divided by (x-1).

OpenStudy (anonymous):

This is really simple. Just learn this simple algorithm. It's an iterative procedure that gives a useful tool for future math. First we find the root of x-1. $$x-1=0 \implies x=1$$ Then we set up a simple figure. See picture. Put the coefficients on top, the root of the divisor on left. Then we drop the first coefficient down. mulitply that by the root and bring up. addup and repeat. Eventually we fill up all the slots below the line. Then this corresponds to $$(x-1) \cdot ((1)\cdot x+(-2) \cdot 1) = \text{the last number in row (In marked it with half box)} =0$$